# Galaxy assembly revealed by globular clusters

### 30 Jan 2024 v1

@@ 1 Introduction
GCs are compact objects with ages typically exceeding 10 Gyr. They serve as ideal tracers of ancient mergers, which have largely dispersed their stellar components into the stellar halos of central galaxies. The sample of MW GCs is relatively more complete than that of single stars and streams, thanks to their compactness and high luminosity. This motivates a growing amount of research (e.g., Massari et al.(2019)Massari, Koppelman, & Helmi; Malhan et al.(2022)Malhan, Ibata, Sharma, Famaey, Bellazzini, Carlberg, D’Souza, Yuan, Martin, & Thomas; Belokurov & Kravtsov(2023)) that uses GCs to probe the MW assembly history. However, these studies rely on different subsets of GC characteristics, such as the age–metallicity space, integral of motion (IoM) space, or orbital action space. Without a solid theoretical basis, these studies must depend on previously established knowledge of mergers, which also vary among the studies. Different classification algorithms used in these studies complicate the comparison of the results. Consequently, current classifications of Galactic GCs are inconsistent with one another, as they are based on different combinations of GC characteristics, prior knowledge, and classification methods.
@@ 2.2 Illustris TNG50 simulations
The model galaxies in the TNG50 simulation are detailed in Chen & Gnedin(2024). We have calibrated the best model parameters to be $(p_{2},p_{3},\kappa)=(14,0.5\,{\rm Gyr^{-1}},1.5)$. The two selected galaxies, 523889 and 519311111The galaxy IDs are SubfindID in the TNG50 subfind catalog., are specifically chosen for their similarity to the Milky Way in terms of virial mass, circular velocity, and key assembly characteristics [^2] The GC catalogs for these two galaxies are publicly accessible at https://github.com/ognedin/gc_model_mw.. We repeat the model on these two galaxies with different random seeds to select the realizations that most accurately reflect the observed total mass, mass function, metallicity distribution, radial profile, and velocity dispersion of the MW GC system.
@@ 2.3.1 Model adaptions to FIRE-2
The optimal parameters for the Latte suite are $(p_{2},p_{3},\kappa)=(7,1\,{\rm Gyr^{-1}},1.5)$. Compared to the best values for TNG50, the $p_{2}$ value is lower due to the higher output frequency of the Latte simulations, by a factor of six. Moreover, the more frequent outputs not only resolve more galaxy assembly events but also record tiny and sometimes spurious mass changes of the host galaxy. In other words, both “signal” and “noise” are magnified. Consequently, we need a higher trigger threshold $p_{3}$ to filter out the noise.
@@ 3.1 Direct clustering
  • Balanced Iterative Reducing and Clustering using Hierarchies (BIRCH) is a hierarchical clustering method in a top-down manner. It begins by constructing a clustering feature (CF) tree from the input data, where each data point is treated as a multi-dimensional vector. The CF for a set has three attributes: the number of elements in the set, the linear sum of all vectors, and the square sum of all vectors. A CF tree node has several sub-clusters, each associated with a CF. A new data point is added to the root of the tree by merging with the sub-cluster that has the smallest radius after merging. If the sub-cluster has child nodes, this merging process continues recursively down the tree. However, if the radius of the sub-cluster exceeds a threshold, and the number of sub-clusters is greater than a branching factor, the node divides. In our implementation, we adopt radius threshold $=0.5$ and branching factor $=50$, following the default values suggested by scikit-learn.
@@
Apart from producing high median accuracy, a robust configuration should avoid the unwanted scenario where the method performs well for some galaxies but yields substantially poorer accuracy for the remaining ones. To assess the robustness of the best configurations, we plot in the right panel of Fig. 3 the maximum and minimum accuracy over the sample galaxies for the same sequence of variables. Some configurations yield significantly low minimum accuracy close to 50%, corresponding to a failed classification. To avoid these cases, we additionally impose a lower threshold for the minimum accuracy. We then repeat the same selection of best configurations but requiring the accuracy for any of the sample galaxies to exceed 70%. Still, we find that the best configurations for Agglomerative Clustering and BIRCH produce median accuracy $\approx 89\%$, as presented in Tab. 1. Particularly, the BIRCH method achieves the highest median accuracy at 89.4% with five variables: $r,E,{\rm[Fe/H]},\varepsilon,J_{\phi}$. In Fig. 4, we illustrate the classification in the ${\rm[Fe/H]}$$J_{r}$ space for the four sample galaxies using this configuration. Among them, the highest accuracy exceeds 90% for 523889, while the lowest accuracy still exceeds 80% for 519311. However, most other configurations fail to achieve a higher accuracy for this galaxy, suggesting a notable variation in the efficiency of identifying in-situ vs. ex-situ GCs across different galaxies. In Appendix B, we quantitatively demonstrate such a galaxy-to-galaxy variation by calculating the maximum achievable accuracy for the four galaxies.
@@ 4.1 Two groups
Figure 7: Classification results for progenitors of ex-situ GCs, with $N_{\rm g}=2$ (top row) and 3 (middle row) for all ex-situ clusters, and $N_{\rm g}=2$ for ex-situ clusters excluding the most dominant merger (bottom row). Although we visualize the results in the ${\rm[Fe/H]}$$J_{r}$ space, the classification is performed with more variables. We use ${\rm age},J_{z},{\rm[Fe/H]},J_{\phi},r_{\rm apo},\varepsilon,$ and $E$ with Agglomerative Clustering for the top row, ${\rm[Fe/H]},\varepsilon,r_{\rm apo},E,$ and ${\rm age}$ with Agglomerative Clustering for the middle row, and ${\rm[Fe/H]},J_{r},J_{z},r_{\rm peri},$ and $E$ with K-means for the bottom row. The colors of filled dots represent the true labels, while those of the open circles stand for the predicted progenitors. For reference, we also show the in-situ clusters as faint red dots, and the clusters from the most dominant merger as faint blue dots in the bottom row.
@@ 5 Application to the Milky Way globular clusters
Throughout this section, we compare our classification results with Massari et al.(2019)Massari, Koppelman, & Helmi and Belokurov & Kravtsov(2023). Massari et al.(2019)Massari, Koppelman, & Helmi classified the progenitors of 151 Galactic GCs, in which 146 are included in our sample. They considered mainly the kinematics of GCs, including the apocenter radius, maximum height above the disk, orbital circularity, specific angular momentum along and perpendicular to the rotational axis, and specific energy. Instead of using unsupervised clustering techniques, their classification is based largely on prior knowledge of the progenitor galaxy properties. Their classification is presented as rectangular regions in these property spaces. If a GC is located in the overlapping region of two rectangles, they refer to this GC as a tentative candidate of both progenitors. We refer to a GC “unambiguous” if these authors assigned the GC to only one unique progenitor.
Belokurov & Kravtsov(2023) classified in-situ vs. ex-situ origins for 165 Galactic GCs with measured kinematics using the same Hilker et al.(2019)Hilker, Baumgardt, Sollima, & Bellini catalogue as in our work. They also applied the same gravitational potential to compute the energy of GCs. Therefore, we can directly compare our classification results with theirs in the IoM space, where these authors used an empirical boundary to separate the in-situ and ex-situ populations, see their Eq. (1). This boundary was carefully selected to isolate in-situ and accreted field stars within the range of metallicities where their classification is robust. They showed that using machine learning methods cannot substantially improve the classification accuracy in the $E-L_{z}$ space.
@@
Using the Bayesian method with an accuracy threshold of 68%, we find that one of the groups resembles the Sagittarius dwarf. This group consists of 13 GCs, located in the high-energy and $L_{z}\gtrsim 0$ region in the IoM space. Remarkably, seven of the eight GCs unambiguously assigned to the Sagittarius dwarf by Massari et al.(2019)Massari, Koppelman, & Helmi fall into this group. The one exception is NGC 5824, which was also not among the six most definite Sagittarius GCs suggested by these authors. No GC from other unambiguous progenitors falls into this group. We detail the classification results for individual GCs in Tab. 4, along with the comparison with Massari et al.(2019)Massari, Koppelman, & Helmi.
@@
[…]
@@ 6.1.2 Impact of the most dominant merger
After the merger with 519311-b, the IoM and action distributions for GCs from all progenitors remained relatively stable over the next 6 Gyr, up to the present. The KL divergence for d, e, and f showed little change, with some even increased slightly up to 30%. This stability suggests that energy and actions are reliable tracers of the progenitor’s properties in the absence of major disruptive merger events. This conclusion is valid in the last 10 Gyr of the MW assembly history, which is relatively quiescent. However, the significant perturbations by GS/E had a significant effect on the kinematics of GCs accreted before the merger, making them less distinguishable. This disruptive merger also explains the difficulty in identifying smaller progenitor galaxies beyond GS/E, as discussed in §4.2.
@@ 6.2.4 Dissolution time and merger duration
Our work differs from Trujillo-Gomez et al.(2023)Trujillo-Gomez, Kruijssen, Pfeffer, Reina-Campos, Crain, Bastian, & Cabrera-Ziri as we specifically focus on MW analog galaxies and aim to classify not only the in-situ vs. ex-situ but also the progenitors of ex-situ clusters. We therefore analyze a selected sample of simulated galaxies matching various observational constraints on the MW. We also take advantage of the available 3D positions and 3D velocities of MW GCs. This allows us to compute more kinematic characteristics in addition to the extragalactic observables. Therefore, although seemingly similar, the motivation and the application areas of the two works are different.
@@ 7 Summary
Next, we concentrate on the ex-situ clusters to distinguish GCs from different progenitor satellite galaxies. To identify the last major merger (GS/E analog), we cluster the ex-situ population into two groups representing the GS/E analog and the other GCs (§4.1). This task can be well handled (accuracy $\gtrsim 90\%$) using seven GC properties: ${\rm age},J_{z},{\rm[Fe/H]},J_{\phi},r_{\rm apo},\varepsilon,$ and $E$ (top row of Fig. 7). Compared to the case of identifying in-situ and ex-situ GCs, two more inputs are needed to distinguish the subtly different distributions of the two groups in the multi-dimensional space. In this case, Agglomerative Clustering and BIRCH are still the most successful clustering methods.
We proceed to identify the second most dominant merger by clustering the ex-situ population into three groups, standing for the GS/E analog, the second most dominant merger, and the other GCs (§4.2). However, no variable configuration yields a minimum accuracy $>70\%$ by directly splitting the ex-situ GCs into three groups (middle row of Fig. 7). If we assume that the GS/E analog can be perfectly labeled and sub-divide only the remaining ex-situ GCs into two groups, we can achieve higher accuracy with a minimum $>70\%$ (bottom row of Fig. 7). Separating the ex-situ population into even more groups is unfeasible. By computing the BIC utilizing a cross-validation technique, we quantitatively show that the overfitting problem becomes severe with more groups (Fig. 8).
@@
Finally, we compare our work with Trujillo-Gomez et al.(2023)Trujillo-Gomez, Kruijssen, Pfeffer, Reina-Campos, Crain, Bastian, & Cabrera-Ziri, who developed a supervised classifier for in-situ vs. ex-situ clusters using the E-MOSAICS simulation (§6.3). Our work outperforms their accuracy in the area of identifying in-situ and ex-situ GCs for MW-analogs because we evaluate our algorithms on the galaxy sample carefully selected to match the MW, and we take into account more kinematic variables that are measurable for the MW GCs.
@@ Acknowledgements.
We are grateful to Andrew Wetzel and Pratik Gandhi for providing the merger trees of the FIRE-2 simulations. We thank Eric Bell, Leandro Beraldo e Silva, Monica Valluri, Andrey Kravtsov, and Yuan-Sen Ting for insightful discussions. OG and YC were supported in part by the U.S. National Science Foundation through grant AST-1909063 and by National Aeronautics and Space Administration through contract NAS5-26555 for Space Telescope Science Institute program HST-AR-16614.

### 19 Mar 2024 v2

@@ 1 Introduction
GCs are compact objects with ages typically exceeding 10 Gyr. They serve as ideal tracers of ancient mergers, which have largely dispersed their stellar components into the stellar halos of central galaxies. The sample of MW GCs is relatively more complete than that of single stars and streams, thanks to their compactness and high luminosity. This motivates a growing amount of research (e.g., Massari et al.(2019)Massari, Koppelman, & Helmi; Forbes(2020); Callingham et al.(2022)Callingham, Cautun, Deason, Frenk, Grand, & Marinacci; Malhan et al.(2022)Malhan, Ibata, Sharma, Famaey, Bellazzini, Carlberg, D’Souza, Yuan, Martin, & Thomas; Belokurov & Kravtsov(2024)) employing GCs to investigate the assembly history of the MW. These studies have utilized diverse classification algorithms on different subsets of GC characteristics, such as the age–metallicity space, integral of motion (IoM) space, or orbital action space. They also relied on previously established knowledge of galactic mergers, which vary across studies. Consequently, while these studies have reached a general agreement on distinguishing in-situ from ex-situ clusters, the assignment of ex-situ GCs to specific progenitors remains debated except for the most significant mergers. This uncertainty arises from the reliance on different combinations of GC characteristics, classification methods, and prior knowledge.
@@ 2.1 Recap of the model
The numerical implementation of the model is publicly available at https://github.com/ybillchen/GC_formation_model under the BSD 3-Clause License, which allows redistribution and modification of the source code with moderate limitations.
@@ 2.2 Illustris TNG50 simulations
The model galaxies in the TNG50 simulation are detailed in Chen & Gnedin(2024). We have calibrated the best model parameters to be $(p_{2},p_{3},\kappa)=(14,0.5\,{\rm Gyr^{-1}},1.5)$. The two selected galaxies, 523889 and 519311111The galaxy IDs are SubfindID in the TNG50 subfind catalog., are specifically chosen for their similarity to the Milky Way in terms of virial mass, circular velocity, and key assembly characteristics. We repeat the model on these two galaxies with different random seeds to select the realizations that most accurately reflect the observed total mass, mass function, metallicity distribution, radial profile, and velocity dispersion of the MW GC system.
@@ 2.3.1 Model adaptions to FIRE-2
The optimal parameters for the Latte suite are $(p_{2},p_{3},\kappa)=(7,1\,{\rm Gyr^{-1}},1.5)$. Compared to the best values for TNG50, the $p_{2}$ value is lower due to the higher output frequency of the Latte simulations, by a factor of six. Moreover, the more frequent outputs not only resolve more galaxy assembly events but also record tiny and sometimes spurious mass changes of the host galaxy. In other words, both “signal” and “noise” are magnified. Consequently, we need a higher trigger threshold $p_{3}$ to filter out the noise. We publicly release the two best MW-analogs from FIRE-2 at https://github.com/ognedin/gc_model_mw, where we have also released the two TNG analogs from Chen & Gnedin(2024).
@@ 3.1 Direct clustering
  • Balanced Iterative Reducing and Clustering using Hierarchies (BIRCH) is a hierarchical clustering method in a top-down manner. It begins by constructing a clustering feature (CF) tree from the input data, where each data point is treated as a multi-dimensional vector. The CF for a set has three attributes: the number of elements in the set, the linear sum of all vectors, and the square sum of all vectors. A CF tree node has several sub-clusters, each associated with a CF. A new data point is added to the root of the tree by merging with the sub-cluster that has the smallest radius after merging. If the sub-cluster has child nodes, this merging process continues recursively down the tree. However, if the radius of the sub-cluster exceeds a threshold, and the number of sub-clusters is greater than a branching factor, the node divides. We find that radius threshold $=0.2$ yields the highest classification accuracy. We adopt branching factor $=50$ following the default values suggested by scikit-learn.
@@
Apart from producing high median accuracy, a robust configuration should avoid the unwanted scenario where the method performs well for some galaxies but yields substantially poorer accuracy for the remaining ones. To assess the robustness of the best configurations, we plot in the right panel of Fig. 3 the maximum and minimum accuracy over the sample galaxies for the same sequence of variables. Some configurations yield significantly low minimum accuracy close to 50%, corresponding to a failed classification. To avoid these cases, we additionally impose a lower threshold for the minimum accuracy. We then repeat the same selection of best configurations but requiring the accuracy for any of the sample galaxies to exceed 70%. Still, we find that the best configurations for Agglomerative Clustering and BIRCH produce median accuracy $\approx 89\%$, as presented in Tab. 1. Particularly, the BIRCH method achieves the highest median accuracy at 89.4% with five variables: $r,E,{\rm[Fe/H]},\varepsilon,J_{\phi}$. In Fig. 4, we illustrate the classification in the ${\rm[Fe/H]}$$J_{r}$ space for the four sample galaxies using this configuration. Among them, the highest accuracy exceeds 90% for 523889, while the lowest accuracy still exceeds 80% for 519311. However, most other configurations fail to achieve a higher accuracy for this galaxy, suggesting a notable variation in the efficiency of identifying in-situ vs. ex-situ GCs across different galaxies. In Appendix B, we quantitatively demonstrate such a galaxy-to-galaxy variation by calculating the maximum achievable accuracy for the four galaxies. We notice that the in-situ clusters misclassified as ex-situ by this classification are more metal-rich than the general ex-situ population. Additionally, these clusters are dynamically hotter than the remaining in-situ GCs, characterized by higher $J_{r}$. These characteristics closely correspond to the “Splash” stellar population identified by Belokurov et al.(2020)Belokurov, Sanders, Fattahi, Smith, Deason, Evans, & Grand. The “Splash” is believed to have originated in the MW’s protodisk and was likely heated by the GS/E merger, resulting in kinematics that resemble those of the ex-situ group. The impact of such early significant mergers will be further studied in §6.1.
@@ 4.1 Two groups
Figure 7: Classification results for progenitors of ex-situ GCs, with $N_{\rm g}=2$ (top row) and 3 (middle row) for all ex-situ clusters, and $N_{\rm g}=2$ for ex-situ clusters excluding the most dominant merger (bottom row). Although we visualize the results in the ${\rm[Fe/H]}$$J_{r}$ space, the classification is performed with more variables. We use ${\rm age},J_{z},{\rm[Fe/H]},J_{\phi},r_{\rm apo},\varepsilon,$ and $E$ with Agglomerative Clustering for the top row, ${\rm[Fe/H]},\varepsilon,r_{\rm apo},E,$ and ${\rm age}$ with Agglomerative Clustering for the middle row, and ${\rm[Fe/H]},J_{r},J_{z},r_{\rm peri},$ and $E$ with K-means for the bottom row. The colors of filled dots represent the true labels, while those of the open circles stand for the predicted progenitors. For reference, we also show the clusters from the most dominant merger as faint blue dots in the bottom row.
@@ 5 Application to the Milky Way globular clusters
Throughout this section, we compare our classification results with Massari et al.(2019)Massari, Koppelman, & Helmi and Belokurov & Kravtsov(2024). Massari et al.(2019)Massari, Koppelman, & Helmi classified the progenitors of 151 Galactic GCs, in which 146 are included in our sample. They considered mainly the kinematics of GCs, including the apocenter radius, maximum height above the disk, orbital circularity, specific angular momentum along and perpendicular to the rotational axis, and specific energy. Instead of using unsupervised clustering techniques, their classification is based largely on prior knowledge of the progenitor galaxy properties. For instance, their criteria to identify in-situ clusters are based on a selected GC sample on the in-situ branch of the age–metallicity relation. Their classification is presented as rectangular regions in these property spaces. If a GC is located in the overlapping region of two rectangles, they refer to this GC as a tentative candidate of both progenitors. We refer to a GC “unambiguous” if these authors assigned the GC to only one unique progenitor.
Belokurov & Kravtsov(2024) classified in-situ vs. ex-situ origins for 165 Galactic GCs with measured kinematics using the same Hilker et al.(2019)Hilker, Baumgardt, Sollima, & Bellini catalogue as in our work. They also applied the same gravitational potential to compute the energy of GCs. Therefore, we can directly compare our classification results with theirs in the IoM space, where these authors used an empirical boundary to separate the in-situ and ex-situ populations, see their Eq. (1). This boundary was carefully selected to isolate in-situ and accreted field stars within the range of metallicities where their classification is robust. Although the two GC populations have distinct distributions in other parameter spaces, such as the $\rm[Mg/Fe]$$\rm[Al/Fe]$ and age–metallicity planes, the boundary in the $E-L_{z}$ space provides a more clear separation between the two populations. They showed that using machine learning methods cannot substantially improve the classification accuracy in the $E-L_{z}$ space.
@@
Using the Bayesian method with an accuracy threshold of 68%, we find that one of the groups resembles the Sagittarius dwarf. This group consists of 13 GCs, located in the high-energy and $L_{z}\gtrsim 0$ region in the IoM space. Remarkably, seven of the eight GCs unambiguously assigned to the Sagittarius dwarf by Massari et al.(2019)Massari, Koppelman, & Helmi fall into this group. The one exception is NGC 5824, which was also not among the six most definite Sagittarius GCs suggested by these authors. The exception is consistent with Bellazzini et al.(2020)Bellazzini, Ibata, Malhan, Martin, Famaey, & Thomas, where NGC 5824 fails to meet all three selection criteria by these authors. No GC from other unambiguous progenitors falls into this group. We detail the classification results for individual GCs in Tab. 4, along with the comparison with Massari et al.(2019)Massari, Koppelman, & Helmi.
@@ 6.1.2 Impact of the most dominant merger
After the merger with 519311-b, the IoM and action distributions for GCs from all progenitors remained relatively stable over the next 6 Gyr, up to the present. The KL divergence for d, e, and f showed little change, with some even increased slightly up to 30%. This stability suggests that energy and actions are reliable tracers of the progenitor’s properties in the absence of major disruptive merger events. This conclusion is valid in the last 10 Gyr of the MW assembly history, which is relatively quiescent. However, the significant perturbations by GS/E had a significant effect on the kinematics of GCs accreted before the merger, making them less distinguishable. This disruptive merger also explains the difficulty in identifying smaller progenitor galaxies beyond GS/E, as discussed in §4.2. A real-world example of such progenitors influenced by the GS/E is the Sequoia event, which likely merged with the Galaxy around the same time as or prior to the GS/E (Myeong et al.(2019)Myeong, Vasiliev, Iorio, Evans, & Belokurov; Forbes(2020); Valenzuela et al.(2023)Valenzuela, Remus, McKenzie, & Forbes). Conversely, the Sagittarius dwarf galaxy merged later, preventing its GC kinematics from being perturbed by the GS/E. Consequently, independent studies have consistently classified the GCs from Sagittarius as a distinct high-energy group with prograde motion in the IoM space.
@@ 6.2.4 Dissolution time and merger duration
Our work differs from Trujillo-Gomez et al.(2023)Trujillo-Gomez, Kruijssen, Pfeffer, Reina-Campos, Crain, Bastian, & Cabrera-Ziri as we specifically focus on MW-analog galaxies and aim to classify not only the in-situ vs. ex-situ but also the progenitors of ex-situ clusters. We therefore analyze a selected sample of simulated galaxies matching various observational constraints on the MW. We also take advantage of the available 3D positions and 3D velocities of MW GCs. This allows us to compute more kinematic characteristics in addition to the extragalactic observables. Therefore, although seemingly similar, the motivation and the application areas of the two works are different.
@@ 7 Summary
Next, we concentrate on the ex-situ clusters to distinguish GCs from different progenitor satellite galaxies. To identify the last major merger (GS/E-analog), we cluster the ex-situ population into two groups representing the GS/E-analog and the other GCs (§4.1). This task can be well handled (accuracy $\gtrsim 90\%$) using seven GC properties: ${\rm age},J_{z},{\rm[Fe/H]},J_{\phi},r_{\rm apo},\varepsilon,$ and $E$ (top row of Fig. 7). Compared to the case of identifying in-situ and ex-situ GCs, two more inputs are needed to distinguish the subtly different distributions of the two groups in the multi-dimensional space. In this case, Agglomerative Clustering and BIRCH are still the most successful clustering methods.
We proceed to identify the second most dominant merger by clustering the ex-situ population into three groups, standing for the GS/E-analog, the second most dominant merger, and the other GCs (§4.2). However, no variable configuration yields a minimum accuracy $>70\%$ by directly splitting the ex-situ GCs into three groups (middle row of Fig. 7). If we assume that the GS/E-analog can be perfectly labeled and sub-divide only the remaining ex-situ GCs into two groups, we can achieve higher accuracy with a minimum $>70\%$ (bottom row of Fig. 7). Separating the ex-situ population into even more groups is unfeasible. By computing the BIC utilizing a cross-validation technique, we quantitatively show that the overfitting problem becomes severe with more groups (Fig. 8).
@@
We compare our work with Trujillo-Gomez et al.(2023)Trujillo-Gomez, Kruijssen, Pfeffer, Reina-Campos, Crain, Bastian, & Cabrera-Ziri, who developed a supervised classifier for in-situ vs. ex-situ clusters using the E-MOSAICS simulation (§6.3). Our work outperforms their accuracy in the area of identifying in-situ and ex-situ GCs for MW-analogs because we evaluate our algorithms on the galaxy sample carefully selected to match the MW, and we take into account more kinematic variables that are measurable for the MW GCs.
Finally, we make the source code, raw data, and key results available in public repositories. The source code for the GC formation model can be accessed at https://github.com/ybillchen/GC_formation_model. The model catalog of GCs is published at https://github.com/ognedin/gc_model_mw. Additionally, the ASCII version of Tab. 4 summarizing our classification of MW GCs is accessible at https://umich.edu/~ognedin/mw_gc_classification.txt.
@@ Acknowledgements.
We are grateful to Andrew Wetzel and Pratik Gandhi for providing the merger trees of the FIRE-2 simulations. We thank Eric Bell, Leandro Beraldo e Silva, Monica Valluri, Andrey Kravtsov, and Yuan-Sen Ting for insightful discussions. We thank the reviewer for their suggestions, which have improved the quality of this work. OG and YC were supported in part by the U.S. National Science Foundation through grant AST-1909063 and by National Aeronautics and Space Administration through contract NAS5-26555 for Space Telescope Science Institute program HST-AR-16614.